OpenStudy (yacoub1993):

Solve and graph the inequality. 4/5x + 5 < -3 A. x < –10

OpenStudy (yacoub1993):

B. x ≤ 0 C. x > 5 D. x ≥ –10

OpenStudy (yacoub1993):

THIS IS CHOICE A GRAPH A. x < –10

OpenStudy (yacoub1993):

CHOICE B GRAPH B. x ≤ 0

OpenStudy (yacoub1993):

CHOICE C GRAPH C. x > 5

OpenStudy (yacoub1993):

CHOICE D GRAPH D. x ≥ –10

OpenStudy (saifoo.khan):

So you tried solving?

OpenStudy (yacoub1993):

YUP i got x=-6.5<-3

OpenStudy (yacoub1993):

is that correct

OpenStudy (saifoo.khan):

No.

OpenStudy (saifoo.khan):

Show me your steps.

OpenStudy (yacoub1993):

ok

OpenStudy (yacoub1993):

\[\frac{ 4 }{ 5 }x+5<-3\] frist i solved the frist equation \[\frac{ 4 }{ 5 }x+5=0\] \[\frac{ 4 }{ 5 }x=-5\] \[x=\frac{ -5 }{ \frac{ 4 }{ 5 } }\] \[x=-6.25\] \[-6.25<-3\] thats the way i did it] correct me if i am wrong

OpenStudy (saifoo.khan):

Hold on a sec. Let me draw.

OpenStudy (yacoub1993):

\[x=-10\] Is that correct

OpenStudy (yacoub1993):

on the graph it will be \[x<-10\]

OpenStudy (saifoo.khan):

Yes sir.

OpenStudy (yacoub1993):

Thanks alot i have another question if u dont mind

OpenStudy (saifoo.khan):

Sure

OpenStudy (yacoub1993):

similar question is it OK

OpenStudy (saifoo.khan):

Yes.

OpenStudy (yacoub1993):

\[\frac{ -x }{ 4 }-6\ge-8\]

OpenStudy (yacoub1993):

A. x < –4 B. x ≤ 8 C. x > 8 D. x ≥ –8

OpenStudy (yacoub1993):

i got \[-x \ge-56\]

OpenStudy (yacoub1993):

wait i will resolve it

OpenStudy (yacoub1993):

Yes i got the same answer bro

OpenStudy (yacoub1993):

i got −x≥−56

OpenStudy (yacoub1993):

@saifoo.khan R u there

OpenStudy (yacoub1993):

@mikala1 Can u help Please

OpenStudy (saifoo.khan):

Sorry i forgot about helping you. lol.

OpenStudy (saifoo.khan):

Your answer's wrong.

OpenStudy (yacoub1993):

what do u think the correct answer?

OpenStudy (yacoub1993):

\[-x \ge-8\]

OpenStudy (saifoo.khan):

Right. But we can't have a negative sign with the variable, right?

OpenStudy (yacoub1993):

what do u mean

OpenStudy (saifoo.khan):

We can NEVER write something with a negative variable. Like here "-x" Negative x is something which we can't write.

OpenStudy (saifoo.khan):

So to get rid of it, we have to flip the inequatlity. I'll show you how..

OpenStudy (saifoo.khan):